# existence1.bib

### v1

@@ 4 Main Results
In this section, we focus on the optimization problem ($P$) and present our main results, including a necessary condition for the finiteness of the optimal value of the problem ($P$) and sufficient conditions for the existence of solutions.
4.1 Necessary Condition for Finiteness of Optimal Value
We provide a result regarding the finiteness of the optimal value of the problem ($P$). To the best of our knowledge, it appears to be new.
Proposition 4.1.1.
Let $X$ be a nonempty set and let $f$ be proper function with $X\cap{\rm dom}f\neq\emptyset$. If the problem ($P$) has a finite optimal value, then
$(f+\delta_{X})_{\infty}(d)\geq 0\qquad\hbox{for all }d\in\mathbb{R}^{n},$
where $\delta_{X}$ is the characteristic function of the set $X$, i.e., $\delta_{X}(x)=0$ when $x\in X$ and $\delta_{X}(x)=+\infty$ otherwise. Moreover, if $X$ and $f$ are additionally assumed to be convex, then we have
$f_{\infty}(d)\geq 0\qquad\hbox{for all }d\in X_{\infty}.$
@@ 4.1 Necessary Condition for Finiteness of Optimal Value
According to the alternative representation of an asymptotic function in Theorem 2.3.1, we have that for any $d\in\mathbb{R}^{n},$
$(f+\delta)_{\infty}(d)=\liminf_{\begin{subarray}{c}d^{\prime}\rightarrow d\\ t\rightarrow+\infty\end{subarray}}\frac{(f+\delta_{X})(td^{\prime})}{t}$
Since $\delta_{X}(td^{\prime})=+\infty$ when $td^{\prime}\notin X$, it follows that
$(f+\delta)_{\infty}(d)=\liminf_{\begin{subarray}{c}td^{\prime}\in X,\,d^{\prime}\rightarrow d\\ t\rightarrow+\infty\end{subarray}}\frac{(f+\delta_{X})(td^{\prime})}{t}\geq\liminf_{t\rightarrow+\infty}\frac{(\inf_{x\in X}f(x))}{t}=0,$
where the inequality follows by using $\delta_{X}(td^{\prime})=0$ and $f(td^{\prime})\geq\inf_{x\in x}f(x)$ when $td^{\prime}\in X$.
When $X$ and $f$ are additionally assumed to be convex, we have
$(f+\delta_{X})_{\infty}(d)=f_{\infty}(d)+(\delta_{X})_{\infty}(d)$
by Remark 3.4.3 in [auslender_teboulle_2002]. Since $(\delta_{X})_{\infty}=\delta_{X_{\infty}}$ by Corollary 2.5.1 of [auslender_teboulle_2002], it follows that
$f_{\infty}(d)+\delta_{X_{\infty}}(d)\geq 0\qquad\hbox{for all }d\in\mathbb{R}^{n},$
implying that $f_{\infty}(d)\geq 0$ for all $d\in X_{\infty}$.
We note that the condition in Proposition 4.1.1 is not sufficient for the finiteness of the optimal value $\inf_{x\in X}f(x)$ even for a convex problem, as seen in the following example.
Example 4.1.
Consider the problem $\inf_{x\in\mathbb{R}}f(x)$ with the function $f(x)=-\sqrt{x}$ for $x\geq 0$, and $f(x)=+\infty$ otherwise. We have that $({\rm epi}f)_{\infty}=\mathbb{R}^{2}_{+}$ and, thus, $f_{\infty}(d)=0$ for all $d\geq 0$, and $f_{\infty}(d)=+\infty$ otherwise. Hence, $f+\delta_{X}\equiv f$ since $X=\mathbb{R}$, and
$(f+\delta_{X})_{\infty}(d)=f_{\infty}(d)\geq 0\qquad\hbox{for all }d\in X_{\infty},$
showing that the condition of Proposition 4.1.1 is satisfied. However, the optimal value of the problem is $f^{*}=-\infty$. ∎
4.2 Conditions for Existence of Solutions to General non-convex Problems
In this section, we provide some sufficient conditions for the existence of solutions for the problem ($P$). The first result relies on the condition that the asymptotic cone $X_{\infty}$ of the set and the asymptotic cone $\mathcal{K}(f)$ of the function have no nonzero vector in common, as given in the following proposition.
Theorem 4.2.1.
Let $X$ be a closed set and $f$ be a proper closed function with $X\cap{\rm dom}f\neq\emptyset$. Assume that $X_{\infty}\cap{\cal K}(f)=\{0\}.$ Then, the problem ($P$) has a finite optimal value $f^{*}$ and its solution set $X^{*}$ is nonempty and compact.
@@ 4.2 Conditions for Existence of Solutions to General non-convex Problems
Since $X\cap{\rm dom}f\neq\emptyset$, there exist a point $x_{0}\in X\cap{\rm dom}f$ with a finite value $f(x_{0})$. Therefore, for $\gamma=f(x_{0})$, the lower-level set $L_{\gamma}(f)$ is nonempty. By Proposition 2.3.1, we have that $(L_{\gamma}(f))_{\infty}\subseteq\mathcal{K}(f)$, thus implying that
where the first inclusion follows from Proposition 2.2.1. Thus, $(X\cap L_{\gamma}(f))_{\infty}=\{0\}$ and the set $X\cap L_{\gamma}(f)$ is bounded by Proposition 2.1.2 of [auslender_teboulle_2002], and hence compact since $X$ and $f$ are closed. Therefore, the problem $\inf_{X\cap L_{\gamma}(f)}f(x)$ has a finite optimal value and a solution exists by the Weierstrass Theorem. Since the problem $\inf_{x\in X\cap L_{\gamma}(f)}f(x)$ is equivalent to the problem ($P$), it follows that $f^{*}$ is finite and attained. The compactness of $X^{*}$ follows by noting that $X^{*}=X\cap L_{f^{*}}(f)$ is nonempty, closed, and bounded due to $(X\cap L_{f^{*}}(f))_{\infty}=\{0\}.$ ∎∎
For the condition on the asymptotic cones of $X$ and $f$ in Theorem 4.2.1, we note that
Theorem 4.2.1 generalizes Theorem 3.1 of [Hieu_2021], which additionally requires that $f$ is bounded below on $X$.
We now focus on a more general case where the intersection $X_{\infty}\cap\mathcal{K}(f)$ contains nonzero directions. We have the following result for the case when $f$ exhibits asymptotically bounded decay with respect to $g(x)=\|x\|^{p}$ on the set $X$, for some $p\geq 0$.
Theorem 4.2.2.
Let the set $X\subseteq\mathbb{R}^{n}$ be closed, and let the objective function $f:\mathbb{R}^{n}\rightarrow\mathbb{R}\cup\{+\infty\}~{}$be proper and closed with $X\cap{\rm dom}f\neq\emptyset$. Assume that $f$ exhibits asymptotically bounded decay with respect to $g(x)=\|x\|^{p}$ on $X$ for some $p\geq 0$, and assume that
Then, the problem ($P$) has a finite optimal value $f^{*}$ and its solution set $X^{*}$ is nonempty.
Proof.
The proof is organized along three steps. The first is to show that a regularized problem of the form $\inf_{x\in X}\{f(x)+r\|x\|^{p}\}$ is coercive and, thus, the regularized problem has a solution for every $r>0$. The second step is to show that any sequence of solutions to the regularized problems, as we vary $r$, is bounded and, thus, has an accumulation point. Finally, we show that every limit point of such a sequence is a solution to the problem ($P$).
@@
By our assumption that $f$ exhibits asymptotically bounded decay with respect to $g(x)=\|x\|^{p}$ on $X$ for some $p\geq 0$, we have that (see Definition 3.1.1)
If $c=+\infty$, then $f$ is coercive on $X$ so by Proposition 2.3.2, the problem ($P$) has a finite optimal value and it has an optimal solution.
Next, we consider the case $c\in\mathbb{R}$.
implying that
Thus, for any $\rho>0$, the regularized function $f(x)+r\|x\|^{p+1}$ is coercive, so by Proposition 2.3.2, the regularized problem
$\inf_{x\in X}\{f(x)+r\|x\|^{p+1}\}$
has a finite optimal value and it has a solution for every $r>0$.
(Step 2: A sequence of solutions to regularized problems is bounded.) Now consider a sequence of positive scalars $\{r_{k}\}$ such that $r_{k}\rightarrow 0$ as $k\rightarrow\infty$. For each $k$, let $x_{k}^{*}\in X$ be a solution to the regularized problem
$\inf_{x\in X}\{f(x)+r_{k}\|x\|^{p+1}\}.$ (9)
We claim that under the conditions of the theorem, the sequence $\{x_{k}^{*}\}\subset X$ must be bounded. To prove this, we argue by contradiction. Assume that $\{x_{k}^{*}\}$ is unbounded. Without loss of generality, we may assume that $x_{k}^{*}\neq 0$ for all $k$, so that the sequence $\{x_{k}^{*}\cdot\|x_{k}^{*}\|^{-1}\}$ is bounded. Hence, it must have a convergent sub-sequence. Without loss of generality, we let $\{x_{k}^{*}\cdot\|x_{k}^{*}\|^{-1}\}\rightarrow d$. Since $\{x_{k}^{*}\}\subset X$, it follows that $d\in X_{\infty}$. Fixing an arbitrary $x_{0}\in X$ we have
$f(x_{k}^{*})\leq f(x_{k}^{*})+r_{k}\|x_{k}^{*}\|^{p+1}\leq f(x_{0})+r_{k}\|x_{0}\|^{p+1}\qquad\hbox{for all $k$,}$
which implies
$\liminf_{k\rightarrow\infty}\frac{f(x_{k}^{*})}{\|x_{k}^{*}\|}\leq\lim_{k\to\infty}\frac{f(x_{0})+r_{k}\|x_{0}\|^{p+1}}{\|x_{k}^{*}\|}=0,$
where in the last equality we use the fact $r_{k}\to 0$. Using the explicit form for the asymptotic function $f_{\infty}$ as given in Theorem 2.3.1, we obtain
Thus, it follows that $d\in X_{\infty}\cap\mathcal{K}(f)$. By our assumption that $X_{\infty}\cap\mathcal{K}(f)\subseteq\mathcal{R}(X)\cap\mathcal{R}(f)$, we have that $d\in\mathcal{R}(X)\cap\mathcal{R}(f)$, that is, $d$ is a retractive direction for the set $X$ and a retractive direction for the function $f$. By the definition of such directions, for the sequence $\{x_{k}^{*}\cdot\|x_{k}^{*}\|^{-1}\}$ and any $\rho>0$, there exists a large enough index $K$ such that
$x_{k}^{*}-\rho d\in X,\qquad f(x_{k}^{*}-\rho d)\leq f(x_{k}^{*})\qquad\hbox{ for all $k\geq K$}.$
Thus, for any $k\geq K$, we have
$f(x_{k}^{*})+r_{k}\|x_{k}^{*}\|^{p+1}\leq f(x_{k}^{*}-\rho d)+r_{k}\|x_{k}^{*}-\rho d\|^{p+1}\leq f(x_{k}^{*})+r_{k}\|x_{k}^{*}-\rho d\|^{p+1}.$
The first inequality follows from the optimality of the point $x_{k}^{*}$ for the regularized problem (9) and the fact that $x_{k}-\rho d\in X$. The second inequality follows by $f(x_{k}^{*}-\rho d)\leq f(x_{k}^{*})$. Since $r_{k}>0$ it follows that for $k\geq K$,
$\|x_{k}^{*}\|^{p+1}\leq\|x_{k}^{*}-\rho d\|^{p+1},$
and therefore,
$\|x_{k}^{*}\|^{2}\leq\|x_{k}^{*}-\rho d\|^{2}\qquad\hbox{for all $k\geq K$}.$
On the other hand, since $\lim_{k\to\infty}x_{k}^{*}\cdot\|x_{k}^{*}\|^{-1}=d$ and $\|d\|=1$, we obtain that $1\leq 0$, which is a contradiction. Hence, it must be that $\{x_{k}^{*}\}$ is bounded.
(Step 3: Any accumulation point of sequence $\{x_{k}^{*}\}$ is a solution to ($P$).) Since $\{x_{k}^{*}\}$ is bounded, it must have an accumulation point. Now, without loss of generality, let $\{x_{k}^{*}\}$ converge to $x^{*}$. Since $x_{k}^{*}\to x^{*}$ and $r_{k}\to 0$, we have that $r_{k}\|x_{k}^{*}\|^{p+1}\to 0$, as $k\to\infty$, implying that
$\liminf_{k\rightarrow\infty}f(x_{k}^{*})=\liminf_{k\rightarrow\infty}\left(f(x_{k}^{*})+r_{k}\|x_{k}^{*}\|^{p+1}\right).$ (10)
Due to the optimality of $x_{k}^{*}$ for the regularized problem, we have for any $x\in X,$
$f(x_{k}^{*})+r_{k}\|x_{k}^{*}\|^{p+1}\leq f(x)+r_{k}\|x\|^{p+1}\qquad\hbox{forall $k$}.$
Therefore, since $r_{k}\to 0$, for any $x\in X$,
$\liminf_{k\rightarrow\infty}\left(f(x_{k}^{*})+r_{k}\|x_{k}^{*}\|^{p+1}\right)\leq f(x)+\lim_{k\rightarrow\infty}r_{k}\|x\|^{p+1}=f(x).$ (11)
Finally, since $x_{k}^{*}\to x^{*}$ and the function $f$ is closed, from relations (10) and (11), it follows that
$f(x^{*})\leq\liminf_{k\rightarrow\infty}f(x_{k}^{*})\leq f(x)\quad\text{ for all }x\in X.$
Since $\{x_{k}^{*}\}\subset X$ and $X$ is closed, we have that $x^{*}\in X$ and $x^{*}$ is an optimal solution to the problem ($P$). ∎∎
The proof of Theorem 4.2.2 is motivated by the proof analysis of Theorem 3.4.1 in [auslender_teboulle_2002], which provides necessary and sufficient conditions for the existence of solutions to an unconstrained minimization problem. Unlike Theorem 3.4.1 in [auslender_teboulle_2002], we consider a more general constrained problem. However, when applied to an unconstrained minimization problem, our Theorem 4.2.2 is weaker than Theorem 3.4.1, since it establishes only sufficient conditions.
Related results for the existence of solutions to $\inf_{x\in X}f(x)$ have been reported in Proposition 12 and Proposition 13 of [bertsekas_tseng_2006] for a convex and a non-convex function $f$, respectively. These results, however, have more stringent requirements than that of Theorem 4.2.2, as discussed later on in Section 6.1. Related is also Proposition 3.1 of [ozdaglar_tseng_2006], which is also discussed in Section 6.1.
Next, we provide an example where the condition $X_{\infty}\cap\mathcal{K}(F)\subseteq\mathcal{R}(X)\cap\mathcal{R}(f)$ is violated and the problem does not have a finite optimal value.
Example 4.2.
Consider the problem of minimizing a convex scalar function $f(x)=-\sqrt{x}$ over its domain $X=\{x\mid x\geq 0\}$. The optimal value is $f^{*}=-\infty$ and there is no solution. Since the function is convex, by Example 3.2, it exhibits asymptotically bounded decay with respect to $g(x)=\|x\|$ on $X$. The set $X$ is a closed convex cone, and we have $X_{\infty}=X$. The asymptotic cone of $f$ coincides with $X$ (see Example 4.1), i.e., $\mathcal{K}(f)=X$, implying that
$X_{\infty}\cap\mathcal{K}(f)=X.$
The cone of retractive directions of $X$ coincides with $X$, since $X$ is a polyhedral set (see Example 3.7). The cone $\mathcal{R}(f)$ of retractive directions of $f$ contains only the zero vector. To see this note that for $x_{k}=\lambda k$, with $\lambda>0$, and $t_{k}=k$ for all $k\geq 1$, we have that $\{x_{k}\}$ converges in the direction $d=\lambda$. However, for any $\rho\in(0,1)$ and any $k\geq 1$,
$f(x_{k}-\rho\lambda)=-\sqrt{k-\rho\lambda}>-\sqrt{k}=f(x_{k}).$
Hence, we have $\mathcal{R}(f)=\{0\}$, and
$\mathcal{R}(X)\cap\mathcal{R}(f)=\{0\},$
thus implying that the condition $X_{\infty}\cap\mathcal{K}(f)\subseteq\mathcal{R}(X)\cap\mathcal{R}(f)$ of Theorem 4.2.2 is violated. ∎
Example 4.3.
Consider the problem of minimizing a convex scalar function $f(x)=e^{x}$ over its domain $X=\mathbb{R}$. The optimal value is $f^{*}=0$ and there is no solution. The function is convex, so by Example 3.2, it exhibits asymptotically bounded decay with respect to $g(x)=\|x\|$ on $X$. We have $X_{\infty}=X$ and $\mathcal{R}(X)=X$. The asymptotic cone of $f$ coincides is given by
$\mathcal{K}(f)=\{x\mid x\leq 0\},$
$f(x_{k}+\rho\lambda)=e^{k+\rho\lambda}>e^{k}=f(x_{k}).$
Hence, we have $\mathcal{R}(f)=\{0\}$ and $\mathcal{R}(X)\cap\mathcal{R}(f)=\{0\},$ implying that the condition $X_{\infty}\cap\mathcal{K}(f)\subseteq\mathcal{R}(X)\cap\mathcal{R}(f)$ of Theorem 4.2.2 is violated. ∎
In Theorem 4.2.2, we used $g(x)=\|x\|^{p}$ for $p\geq 0$. We now show the existence result in the case of a coercive function $g$.
Theorem 4.2.3.
Let the feasible set $X\subseteq\mathbb{R}^{n}$ be closed, and let the objective function $f:\mathbb{R}^{n}\rightarrow\mathbb{R}\cup\{+\infty\}~{}$be proper and closed. Suppose that $f$ exhibits asymptotically bounded decay with respect to a coercive function $g$ on $X$ and that $X\cap{\rm dom}f\cap{\rm dom}g\neq\emptyset$. Assume further that
Let $g$ have Lipschitz continuous gradients on $\mathbb{R}^{n}$ and, for any sequence $\{x_{k}\}\subseteq X$ converging in any nonzero direction $d\in\mathcal{R}(X)\cap\mathcal{R}(f)$, let the following relation hold
$\limsup_{k\to\infty}\frac{\langle\nabla g(x_{k}),d\rangle}{t_{k}}>0\qquad\hbox{where }\lim_{k\to\infty}\frac{x_{k}}{t_{k}}=d.$
Then, the problem ($P$) has a finite optimal value and its solution set is nonempty.
Proof.
Consider the regularized problem for any $r>0$,
$\inf_{x\in X}\{f(x)+rg^{2}(x)\}.$ (12)
By our assumption that $f$ exhibits asymptotically bounded decay with respect to $g$ on $X$, we have by Definition 3.1.1 that
@@
If $c=+\infty$, then $f$ is coercive since $g$ is coercive, and the result follows by Proposition 2.3.2.
Next, we consider the case $c\in\mathbb{R}$.
@@
where the equality is due to $r>0$ and the coercivity of $g$ on $X$. Hence, the regularized objective function itself is also coercive, and by Proposition 2.3.2 the problem $\inf_{x\in X}\{f(x)+rg^{2}(x)\}$ has a solution for any $r>0$.
@@
where we use the fact $r_{k}\to 0$. Employing Theorem 2.3.1 which gives the explicit form for the asymptotic function $f_{\infty}$, we can see that
@@
$g^{2}(x_{k}^{*})\leq g^{2}(x_{k}^{*}-\rho d)\qquad\hbox{for all $k\geq K$}.$
Assume that $K$ is large enough so that $g(x_{k})>0$ for all $k\geq K$. Then, the preceding relation implies that
$0<g(x_{k}^{*})\leq g(x_{k}^{*}-\rho d)\leq g(x_{k}^{*})-\rho\langle\nabla g(x_{k}^{*}),d\rangle+\frac{\rho^{2}}{2L}\qquad\text{ for all }k\geq K.$ (16)
where in the last inequality we use the Lipshitz continuity of $\nabla g$ and $\|d\|=1$. Relation (16) and the fact that $\rho>0$ imply that
$\langle\nabla g(x_{k}^{*}),d\rangle\leq\frac{\rho}{2L}\qquad\text{ for all }k\geq K.$
Hence, it follows that
$\limsup_{k\to\infty}\frac{\langle\nabla g(x_{k}^{*}),d\rangle}{\|x_{k}^{*}\|}\leq\limsup_{k\to\infty}\frac{\rho}{2L\|x_{k}^{*}\|}=0,$
which contradicts the assumption on $\nabla g$ that $\limsup_{k\to\infty}\frac{\langle\nabla g(x_{k}^{*}),d\rangle}{\|x_{k}^{*}\|}>0$. Thus, the sequence $\{x_{k}^{*}\}$ must be bounded.
@@
Since $\{x_{k}^{*}\}\subseteq X$ and the set $X$ closed, we have that $x^{*}\in X$ is an optimal solution to the problem ($P$). ∎∎
We have the following result as a special consequence of Theorem 4.2.2 (or Theorem 4.2.3) for the case when the set $X$ is polyhedral.
Corollary 4.2.1.
Let $X$ be a nonempty polyhedral set and $f$ be a proper closed function with $X\cap{\rm dom}f\neq\emptyset$. Under assumptions of Theorem 4.2.2 (or Theorem 4.2.3) on the asymptotically bounded decay of $f$ on $X$, if
$X_{\infty}\cap\mathcal{K}(f)\subseteq X_{\infty}\cap\mathcal{R}(f),$
then the problem ($P$) has a finite optimal value and a solution exists.

### v3

@@ 4 Main Results
In this section, we consider the following minimization problem
$\inf_{x\in X}f(x),$ ($P$)
where $X\subseteq\mathbb{R}^{n}$ is a nonempty closed set and $f$ is a proper closed function. We let $f^{*}$ denote the optimal value of the problem, and $X^{*}$ denote the set of optimal solutions. We provide our sufficient conditions for the existence of solution results in Subsection 4.1 and relate them to some of the existing conditions in Subsection 4.2.
4.1 Sufficient Conditions for Existence of Solutions
The first result relies on the condition that the asymptotic cone $X_{\infty}$ of the set and the asymptotic cone $\mathcal{K}(f)$ of the function have no nonzero vector in common.
Theorem 4.1.
Let $X$ be a closed set and $f$ be a proper closed function with $X\cap{\rm dom}f\neq\emptyset$. Assume that $X_{\infty}\cap{\cal K}(f)=\{0\}.$ Then, the problem ($P$) has a finite optimal value, and nonempty and compact solution set $X^{*}$.
@@ 4.1 Sufficient Conditions for Existence of Solutions
Since $X\cap{\rm dom}f\neq\emptyset$, there exists a point $x_{0}\in X\cap{\rm dom}f$ with a finite value $f(x_{0})$. Therefore, for $\gamma=f(x_{0})$, the lower-level set $L_{\gamma}(f)$ is nonempty. By Proposition 2.3, we have that $(L_{\gamma}(f))_{\infty}\subseteq\mathcal{K}(f)$, thus implying that
where the first inclusion follows from Proposition 2.1. Thus, $(X\cap L_{\gamma}(f))_{\infty}=\{0\}$ and the set $X\cap L_{\gamma}(f)$ is bounded by Proposition 2.1.2 of [2], and hence compact since $X$ and $f$ are closed. Therefore, the problem $\inf_{X\cap L_{\gamma}(f)}f(x)$ has a finite optimal value and a solution exists by the Weierstrass Theorem. Since the problem $\inf_{x\in X\cap L_{\gamma}(f)}f(x)$ is equivalent to the problem ($P$), it follows that $f^{*}$ is finite and attained. The compactness of $X^{*}$ follows by noting that $X^{*}=X\cap L_{f^{*}}(f)$ is nonempty, closed, and bounded due to $(X\cap L_{f^{*}}(f))_{\infty}=\{0\}.$
For the condition on the asymptotic cones of $X$ and $f$ in Theorem 4.1, we note that
Theorem 4.1 generalizes Theorem 3.1 of [12] for minimizing a polynomial function, which requires that $f$ is bounded below on $X$. It is also more general than Theorem 3 of [10], when restricted to a finite dimensional norm space, which requires that $X$ is convex and $f$ is bounded below on $X$. However, Theorem 3 of [10] applies to problems in infinite-dimensional spaces and, by virtue of this, it is more general than our Theorem 4.1.
We make a final remark that the condition of Theorem 4.1 implies that
$(f+\delta_{X})_{\infty}(d)>0\qquad\hbox{for all nonzero $d\in\mathbb{R}^{n}$},$
where $\delta_{X}$ is the indicator function of the set $X$.
Thus, the function $f+\delta_{X}$ satisfies the conditions of Theorem 3.4.1 in [2], which implies that the problem ($P$) has a solution, but it does not imply that the solution set is bounded.
We now provide sufficient conditions for the existence result using the asymptotically bounded decay property with respect to a coercive function $g$.
Theorem 4.2.
Let the set $X\subseteq\mathbb{R}^{n}$ be closed, and let the function $f:\mathbb{R}^{n}\rightarrow\mathbb{R}\cup\{+\infty\}~{}$be proper closed and and $X\cap{\rm dom}f\neq\emptyset$. Suppose that $f$ exhibits asymptotically bounded decay with respect to a coercive function $g$ on $X$. Assume that
Let $g$ have Lipschitz continuous gradients on $X$ and, for any sequence $\{x_{k}\}\subseteq X$ converging in any nonzero direction $d\in\mathcal{R}(X)\cap\mathcal{R}(f)$, let the following relation hold
$\liminf_{k\to\infty}\frac{\langle\nabla g(x_{k}),d\rangle}{t_{k}}>0\qquad\hbox{where }\lim_{k\to\infty}\frac{x_{k}}{t_{k}}=d.$
Then, the problem ($P$) has a finite optimal value $f^{*}$ and a nonempty solution set $X^{*}$.
@@
By our assumption that $f$ exhibits asymptotically bounded decay with respect to $g$ on $X$, we have by Definition 3.1 that
@@
If $c=+\infty$, then $f$ is coercive since $g$ is such. Thus, the result follows by Proposition 2.4.
Next, we consider the case $c\in\mathbb{R}$. For this, we consider the regularized problem
$\inf_{x\in X}\{f(x)+rg^{2}(x)\}\qquad\hbox{for any $r>0$}.$ (9)
@@
where the equality is due to $r>0$ and the coercivity of $g$ on $X$. Hence, the regularized objective function itself is also coercive, and by Proposition 2.4 the problem $\inf_{x\in X}\{f(x)+rg^{2}(x)\}$ has a solution for any $r>0$.
@@
where we use the fact $r_{k}\to 0$. Employing Theorem 2.2 which gives the explicit form for the asymptotic function $f_{\infty}$, we can see that
@@
$g^{2}(x_{k}^{*})\leq g^{2}(x_{k}^{*}-\rho d)\qquad\hbox{for all $k\geq K$}.$ (12)
Assume that $K$ is large enough so that $g(x_{k})>0$ for all $k\geq K$. Then, relation (12) implies that
$0<g(x_{k}^{*})\leq g(x_{k}^{*}-\rho d)\leq g(x_{k}^{*})-\rho\langle\nabla g(x_{k}^{*}),d\rangle+\frac{\rho^{2}L}{2}\qquad\text{ for all }k\geq K.$ (13)
where in the last inequality we use the Lipschitz continuity of $\nabla g$ and $\|d\|=1$. Since $\rho>0$, by relation (13) it follows that
$\langle\nabla g(x_{k}^{*}),d\rangle\leq\frac{\rho L}{2}\qquad\text{ for all }k\geq K.$
Hence,
$\liminf_{k\to\infty}\frac{\langle\nabla g(x_{k}^{*}),d\rangle}{\|x_{k}^{*}\|}\leq\liminf_{k\to\infty}\frac{\rho L}{2\|x_{k}^{*}\|}=0,$
which contradicts the assumption on $\nabla g$ that $\liminf_{k\to\infty}\frac{\langle\nabla g(x_{k}^{*}),d\rangle}{\|x_{k}^{*}\|}>0$. Thus, the sequence $\{x_{k}^{*}\}$ must be bounded.
@@
Since $\{x_{k}^{*}\}\subseteq X$ and $X$ closed, $x^{*}\in X$ is an optimal solution to the problem ($P$). ∎
The proof of Theorem 4.2 is motivated by the insights gained from proof of Theorem 3.4.1 in [2], which is in turn motivated by [3], and it has ideas similar to work [17] on the existence of solutions for problems (in a Banach space).
The following result is a consequence of
Theorem 4.2 for a polyhedral set $X$.
Corollary 4.1.
Let $X$ be a polyhedral set and $f$ be a proper closed function with $X\cap{\rm dom}f\neq\emptyset$. Under assumptions of Theorem 4.2, if $X_{\infty}\cap\mathcal{K}(f)\subseteq X_{\infty}\cap\mathcal{R}(f),$ then the problem ($P$) has a solution.